OpenStudy (anonymous):

find the axis of symmetry for this parabola: y=-4x^2+8x-12 Write as equation.

OpenStudy (anonymous):

Axis of symmetry will run through the vertex of this. What we know about the vertex of a parabola is the slope will be 0. I'd personally find the derivative (y'=-8x+8), set y' to 0, solve for x (which is 1), and that'll make x=1 your answer. It should be written as an equation (x=1) because this is a line, and '1' is just a number, not a line.

OpenStudy (anonymous):

y=-4(x^2+8x____)-12 thats all i know

OpenStudy (anonymous):

Ah, I see. You're completing the square to find the vertex. That's possible, too. Don't forget to factor a -4 out of your 8x, too. That'll give you (y+8) and (x-1), which is the point (1,-8). X still equals 1 because this is an axis, and will either be x=something or y=something (it's either straight up or straight sideways). In this case, it's up/down meaning x=something, and your answer is still x=1.

OpenStudy (anonymous):

Oh, and by completing the square, you should have gotten y+8=-4(x-1)^2. GL!

OpenStudy (anonymous):

but what would the equation be for the axis of symetry... x=1

OpenStudy (anonymous):

Yep. Just x=1, which is a simple equation, but still an equation! :) A line straight up is always x=(number) with no y-value.